OpenStudy (anonymous):

How do you find sin A and tan A if cosA equals 4/5 and cot A >0.

OpenStudy (b87lar):

cosA=4/5 corresponds to a right angled triangle with one side=4 and hypotenuse=5 From Pythagoras you then can tell the third side, which is 3. so \[\sin A=\frac{3}{5}\] and \[\tan A = \frac{3}{4}\]

OpenStudy (b87lar):

Because we know that cotA>0 we also know that sinA cannot be negative (while cosA is given to be positive), so we have excluded the possibility that sinA=-3/5 (which otherwise could have been a possible solution, too)

OpenStudy (b87lar):

Does it make sense?

OpenStudy (anonymous):

i dont understand how you know if it is positive or negative like the cot A >0 confuses me

OpenStudy (anonymous):

you must be able to draw a circle first :DDD

OpenStudy (b87lar):

ok. let's look at the definition of cot:\[\cot A = \frac{\sin A}{\cos A} > 0 \] This can be true only if either \[\sin A > 0, \cos A>0\] or \[\sin A <0, \cos A <0\] agree?

OpenStudy (b87lar):

Once you agree with that you can exclude the second possibility, since we're told that cosA=4/5 > 0 --> so sinA cannot be < 0.

OpenStudy (b87lar):

sorry i correct myself cot should be \[\cot A=\frac{\cos A}{\sin A}\] but the argument is the same...

OpenStudy (anonymous):